OpenStudy (marigirl):

What is the mathematical term to describe the a value of a parabola y=ax^2+bx+c What i mean is the steepness of the parabola...what is the mathematical term

ganeshie8 (ganeshie8):

In the polynomial \(ax^2+bx+c\), \(ax^2\) is called the leading term and, \(a\) is called "leading coefficient"

OpenStudy (marigirl):

yes but when using that to describe a parabola such as y=x^2 and y=0.5x^2 when looking at the Domain values [0,5] we can see that y=x^2 graph is more steep than 0.5x^2... it is because of the coefficient..is there a special mathematical term for this? or should i just use plain English.

OpenStudy (welshfella):

I don't know of a mathematical term for this. In this case plain English is, I believe , , perfectly adequate.